OpenStudy (anonymous):

determine which of the following are subspaces of the space F(-infinity,infinity) (a) all f such that f(x)< or equal to 0 for all x (b) all f such that f(0)=0 (c) all f such that f(0)=2 (d) all constant functions (e) all f of the form k1+k2sinx , where k1 and k2 are real numbers

OpenStudy (anonymous):

What's F(-infinity,infinity)?

OpenStudy (anonymous):

F denotes continuous functions

OpenStudy (anonymous):

No D is ok

OpenStudy (anonymous):

please explain sir

OpenStudy (anonymous):

Never mind my answer...

OpenStudy (anonymous):

Add two constant functions, you have a constant function Scalar multiple of a constant function, you have a constant function Zero vector exists

OpenStudy (anonymous):

A) fails scalar mult B) ok C) fails scalar mult D) ok

OpenStudy (anonymous):

what about e

OpenStudy (anonymous):

Seems OK to me, passes all three tests.

OpenStudy (anonymous):

D also fails addition property

OpenStudy (anonymous):

D? you can add two constant functions to get a constant function

OpenStudy (anonymous):

mistake i meant C fails addition property

OpenStudy (anonymous):

is E is a subspace

OpenStudy (anonymous):

is C fails addition property

OpenStudy (anonymous):

yes because suppose you have two such functions, f + g (0) = 4

OpenStudy (anonymous):

what about E

OpenStudy (anonymous):

I think its a subspace, but verify yourself. Test all the axioms.

OpenStudy (anonymous):

i think it is can you check please sir

OpenStudy (anonymous):

k_1 + k_2*sin(x) + k_3 + k_4*sin(x) = (k_1 + k3) + (k_2+ k_4)*sin(x) closure under addition zero vector where k_1= 0 and k_2=0 and scalar mult is fine

OpenStudy (anonymous):

i do not understand zero vector can you please explain

OpenStudy (anonymous):

A subspace must satisfy the following: It must have a zero vector It must be closed under addition It must be closed under multiplication

OpenStudy (anonymous):

If k_1 = 0 and k_2 = 0 then we have 0+ 0*sin(x) = 0